OpenStudy (anonymous):

For the equation x2 + 3x + j = 0, find all the values of j such that the equation has two real number solutions. Show your work

OpenStudy (anonymous):

for 2 real solutions the discriminant b^2 - 4ac > 0 here it = 3^2 - 4 *1*j = 9 - 4j solve the inequality 9 - 4j = 0 to find required values of j

OpenStudy (anonymous):

the a, b and c of the discriminant formula are from the general equation of quadratic ax^2 + bx + c = 0

OpenStudy (anonymous):

ax^2 + bx + c = 0 compare with x^2 + 3x + j = 0 a=1, b = 3 and c = j

OpenStudy (anonymous):

can you solve 9 - 4j > 0 ?